OpenStudy (anonymous):

Josh made $252 for 18 hours of work. At the same rate, how much would he make for 11hours of work?

OpenStudy (anonymous):

@peterg2001 @magepker728

OpenStudy (peterg2001):

252 dived by 18=14 14x11=154 i think the answer is 154

OpenStudy (stamp):

$252 = 18 hours of work Divide on 18 by each side to see how much one hour of work is $252/18 = 1 hour of work Now multiply each side by eleven to solve for 11 hours of work.

OpenStudy (stamp):

Or set up a proportion to solve \[\frac{$252}{18}=\frac{x}{11}\]

OpenStudy (peterg2001):

@stamp am i wrong?

OpenStudy (stamp):

Probably not